use the transformation x=et to transform the equation x2d2y/dx2-2xdy/dx+2y=x3 to a linear equation with constant coefficients hence solve the equation


1
Expert's answer
2021-08-02T14:40:34-0400

Let us solve the differential equation "x^2\frac{d^2y}{dx^2}-2x\frac{dy}{dx}+2y=x^3." Let us use the transformation "x=e^t." Then "y'_x=y'_t\cdot t'_x=y'_t\cdot \frac{1}{x'_t}=y'_t\cdot e^{-t}," "y''_{x^2}=(y''_{t^2}\cdot e^{-t}-y'_t\cdot e^{-t})e^{-t}=(y''_{t^2}-y'_t)e^{-2t}."

It follows that we have the following equation "e^{2t}(y''_{t^2}-y'_t)e^{-2t}-2e^ty'_te^{-t}+2y=e^{3t}," which is equivalent to the equation "y''_{t^2}-3y'_t+2y=e^{3t}." The characteristic equation "k^2-3k+2=0" is equivalent to "(k-1)(k-2)=0," and hence has the roots "k_1=1" and "k_2=2." The general solution of the differential equation "y''_{t^2}-3y'_t+2y=e^{3t}" is "y(t)=C_1e^t+C_2e^{2t}+y_p(t)," where "y_p(t)=ae^{3t}." It follows that "y_p'(t)=3ae^{3t},y_p''(t)=9ae^{3t}." We have that "9ae^{3t}-9ae^{3t}+2ae^{3t}=e^{3t}." Then "2a=1," and thus "a=\frac{1}{2}." Therefore, "y(t)=C_1e^t+C_2e^{2t}+\frac{1}{2}e^{3t}." We conclude that the general solutions of the differential equation "x^2\frac{d^2y}{dx^2}-2x\frac{dy}{dx}+2y=x^3" is "y(x)=C_1x+C_2x^2+\frac{1}{2}x^3."